Spectrum Protect (TSM): maintenance affecting restores from backup-L and oc11-bk-ent-1
Scheduled Maintenance Report for MIT
Completed
The scheduled maintenance has been completed.
Posted Jun 15, 2022 - 16:00 EDT
In progress
Scheduled maintenance is currently in progress. We will provide updates as necessary.
Posted Jun 15, 2022 - 14:00 EDT
Scheduled
The fiber optic connections that allow backup-L and oc11-bk-ent-1 to communicate with tape drives will undergo maintenance (a firmware upgrade). If all goes well, the work will take around an hour. However, if there are issues we may need to work with our vendor to resolve them, and that will take longer.
Please note that restore processes on backup-L and oc11-bk-ent-1 may fail if they are reading a tape during certain portions of this work. This work will not affect desktop restores via Code42/Crashplan.
Posted Jun 02, 2022 - 16:38 EDT